# Stationary states of forced two-phase turbulence

### v1

@@ 1 Introduction
The deformation, fragmentation, and coalescence of bubbles and droplets in a turbulent environment is an important problem in various environmental and industrial settings (Liao & Lucas, 2009, 2010; Jain, 2017; Ni, 2024). In the atmosphere, the enhanced efficiency of water droplet collision–coalescence events due to air turbulence heavily influences cloud formation and warm rain initiation (Grabowski & Wang, 2013). At the ocean–atmosphere interface, underwater turbulence generated due to breaking waves breaks apart large entrained air pockets into smaller bubbles, facilitating mass transfer between the ocean and the atmosphere (Deike, 2022). Finally, turbulent emulsions, i.e., a homogenized mixture of two fluids of similar densities, are ubiquitously formed for use in pharmaceutical and industrial processing (Wang et al., 2007; Kilpatrick, 2012). As such, an understanding of how turbulence influences (and is influenced by) the dispersed phase, as parameters such as density and viscosity ratios, between the carrier and dispersed phases, as well as the volume fraction occupied by the dispersed phase are varied, is of paramount importance to understanding these ubiquitous flows.
@@
While these examples contain irreducible effects due to anisotropy, large-scale shear (Rosti et al., 2019), and buoyancy (Ravelet et al., 2011), the most straightforward setup to study the essential physics of deformation, fragmentation, and coalescence of bubbles and droplets in a turbulent environment is homogeneous isotropic turbulence (HIT). While simplistic, this setup allows for systematically studying the high-dimensional parameter space of density and viscosity ratios, volume fraction of dispersed phase, and Weber and Reynolds numbers. For each of these global parameters, the dependence of the phenomenology of the flow on the scales of turbulent fluctuations in the carrier phase and the scales of the dispersed phase can be further probed both above and below the Hinze scale (Hinze, 1955) (scale at which the inertial forces causing fragmentation due to turbulence are balanced by the restorative forces of surface tension). Incorporating the appropriate phenomenology of fragmentation and coalescence as a function of the global and scale-dependent parameters is essential to develop coarse-grained statistical models that can accurately capture the quantitative properties of the mixture. For example, while similar density mixtures lead to large-scale filamentary deformations requiring capillary-like breakup models (Eastwood et al., 2004; Kim & Moin, 2020), large density ratios such as in the case of air-water bubble break-up might be better modeled using a cascade-like model that may or may not account for the intermittent nature of turbulence (Chan et al., 2021a, b; Qi et al., 2022).
@@
Recent studies that utilize two-phase HIT as a testbed can be broken down into two categories: (i) The first category considers the detailed dynamics of individual bubbles and droplets, or emulsions, during the first few breakup events through simulations containing a single bubble/droplet, characterizing and quantifying things such as the generation of sub-Hinze scale bubbles due to rapid successive breakup events and capillary fragmentation (Riviere et al., 2021; Rivière et al., 2022), the effect of eddies far away from the two-phase interface on the deformation process (Vela-Martín & Avila, 2021), bubble deformation modes due to turbulent fluctuations (Perrard et al., 2021), the non-vanishing probability of breakup at sub-Hinze scales due to increasingly intermittent turbulent fluctuations (Vela-Martín & Avila, 2022), and the role of the viscosity ratio between the two phases on the rate of bubble breakup and the critical Weber number at which they break (Farsoiya et al., 2023). (ii) The second category considers characterizing the statistical properties of numerous bubbles and droplets present simultaneously in a turbulent environment. Within this category, Dodd & Ferrante (2016) studied the turbulent kinetic energy (TKE) budgets of droplets at various density ratios within decaying HIT. Crialesi-Esposito et al. (2022, 2023) studied the influence of viscosity ratios and volume fractions of the dispersed phase at unity density ratios in a forced HIT setting on the scale-by-scale energy budgets while linking the droplet-size distribution (DSD) to the local dissipation rate of turbulence, once again highlighting the importance of intermittency in the breakup dynamics. Finally, Begemann et al. (2022) and Krzeczek et al. (2023) utilize a linear forcing approach to study the segregation properties of emulsions, highlighting that even minor variations in the density ratio of the two phases from slightly below to slightly above unity shifts the DSD to have fewer super-Hinze bubbles and more sub-Hinze scale drops.
@@
Beyond this uncertainty of non-physical initial conditions, studies using decaying turbulence cannot separate the effect of the non-equilibrium change in the global properties of the surrounding carrier phase turbulence on the extracted statistics of the fragmentation processes for droplets/bubbles with timescales commensurate with some of the energetic scales of turbulence. In particular, for bubbles/droplets placed in the inertial range of turbulence, the recent evidence for the importance of a range of scales of smaller eddies on the break-up process could be drastically different if those small scales are decaying in time (Qi et al., 2022). Controlling this non-equilibrium effects would allow for more accurate measurements of the equilibrium statistics against which the non-equilibrium cases can be compared and interpreted as either slight deviations or completely different phenomena. As such, a statistically stationary state is sought instead. To achieve this state, forcing is necessary to sustain the continuously dissipating turbulence, ensuring that the turbulence properties do not change during the breakup processes. In the cases where forcing has been used, both simultaneous (both phases) and individual phase (carrier), linear, physical-space forcing (Lundgren, 2003; Rosales & Meneveau, 2005) have been utilized in various studies. However, individual phase (carrier) forcing is only used for the cases that study the break-up of single bubbles in HIT (Riviere et al., 2021; Rivière et al., 2022), and are hence more likely influenced by the choice of the initial condition as the simulations are only performed up to the first break-up event. In contrast, simultaneous mixture-phase (both phases) forcing has been applied only to density ratios representative of emulsions (Begemann et al., 2022; Crialesi-Esposito et al., 2022; Krzeczek et al., 2023; Crialesi-Esposito et al., 2023). As such, there remain open questions concerning the differing effects of simultaneous mixture-phase (both phases) forcing and individual phase (carrier) forcing, as well as how the aggregate properties of the two-phase HIT vary with density and viscosity ratios when forced.
@@ 2.1 Phase-field model
Note that a small number $\varepsilon$ is added to both the numerator and denominator to avoid division by 0, and $\varepsilon$ is chosen as $\varepsilon=10^{-100}$. The right-hand-side (RHS) term of Eq. (1) is the interface-regularization term, which is an artificial term that is added to maintain the constant thickness of the interface. The interface-regularization term contains a diffusion and a sharpening term, which balance each other to maintain a constant interface thickness on the order of $\epsilon$. The ACDI model is discretized using a second-order skew-symmetric split central scheme (Jain & Moin, 2022) that has superior numerical properties, such as improved robustness and lower aliasing error.
@@
This recently developed ACDI method has been shown to be suitable for complex turbulent flow simulations (Jain, 2022; Hwang & Jain, 2024) and a wide range of other multiphysics flow problems (Collis et al., 2022; Scapin et al., 2022; Brown et al., 2023; Jain, 2023, 2024). It is known to be more accurate than other existing phase-field models because it maintains a sharper interface—with only one-to-two grid points across the interface—while being non-dissipative, robust, less expensive, and conservative, without the need for any geometric treatment. Because of these advantages, the ACDI method is chosen as the interface-capturing method in this work.
@@ 2.3 Surface tension model
The surface tension forces, $f^{\sigma}_{i}=\sigma\kappa(\partial\phi/\partial x_{i})$ [last term in Eq. (3)], are modeled using a continuum-surface force (CSF) formulation (Brackbill et al., 1992), where $\sigma$ is the surface tension coefficient and $\kappa$ is the surface curvature, defined as $\kappa=-(\partial n_{j}/\partial x_{j})=-\partial/\partial x_{j}\left[(\partial\psi/\partial x_{j})/\sqrt{(\partial\psi/\partial x_{i})(\partial\psi/\partial x_{i}})\right]$. The use of $\psi$, as opposed to $\phi$, to compute curvature in the surface tension force was shown in Jain (2022) to improve modeling of the surface tension force and reduce spurious currents significantly. This improved CSF model results in lower spurious currents compared to other existing surface tension models for diffuse-interface methods in the literature, including the free-energy based surface tension models. Hence, the improved CSF formulation is used to model surface tension forces in this work.
@@ 3.1 Mixture energy equations
where $k=\langle\rho u_{i}u_{i}/2\rangle$ is the mean mixture TKE, $\Psi=\langle u_{i}f_{i}^{\sigma}\rangle$ is the mean surface tension power, and $\epsilon=\langle\tau_{ij}{\partial u_{i}}/{\partial x_{j}}\rangle$ is the dissipation; for all these terms, the averaging operator is defined as $\langle\cdot\cdot\rangle=\left[\int_{\Omega}(\cdot\cdot)d\Omega\right]/\Omega$, where $\Omega$ is the domain volume. Similarly, we can write the mean surface energy equation (Dodd & Ferrante, 2016) as
@@ 3.2 Phase-averaged energy equations
where $k_{l}=\langle\phi_{l}\rho_{l}u_{i,l}u_{i,l}/2\rangle/\langle\phi_{l}\rangle$ is the phase-averaged TKE, $V_{l}=\langle\phi_{l}\left({\partial u_{i,l}\tau_{ij,l}}/{\partial x_{j}}\right)\rangle/\langle\phi_{l}\rangle$ is the phase-averaged viscous transport contribution, $P_{l}=\langle\phi_{l}\left({\partial u_{j,l}p_{l}}/{\partial x_{j}}\right)\rangle/\langle\phi_{l}\rangle$ is the phase-averaged pressure transport contribution, and $\epsilon_{l}=\langle\phi_{l}\tau_{ij}\left({\partial u_{i,l}}/{\partial x_{j}}\right)\rangle/\langle\phi_{l}\rangle$ is the dissipation of phase $l$. These equations are consistent with the diffuse-interface/phase-field formulation, where $\phi_{l}$ here represents the phase-field variable [see Dodd & Ferrante (2016) for a sharp-interface formulation of phase-averaged TKE equations].
@@ 5 Spatial filtering approach for two-phase spectra
To study the energy content at different scales of each of the dispersed and carrier phases in a two-phase flow, we compute the phase-dependent TKE spectra utilizing the spatial filtering approach of Sadek & Aluie (2018). Note that a standard Fourier spectrum is not a general enough approach for when there are variations in density and viscosity between the two phases. This is due to three reasons: First, the viscosity variations introduce discontinuities of velocity derivatives at interfaces, contaminating the spectrum. Second, using a Fourier spectrum, one cannot isolate the dispersed phase and carrier phase contributions. Third, and perhaps most importantly, energy is a cubic quantity in compressible flows and volumetrically filtered variable-density flows such as non-unity density ratio incompressible two-phase flows. However, energy must be defined quadratically using a Fourier spectrum; no longer satisfying the inviscid criterion (Aluie, 2013; Zhao & Aluie, 2018), i.e., the lack of influence of viscous effects at the large scales.
Using the wavelet method of Freund & Ferrante (2019), spectra have been computed before for the dispersed and carrier phases for some wavenumbers in an incompressible two-phase flow. However, due to the need for decomposing spatial regions into carrier, dispersed, and interaction phases, the range of scales over which a dispersed phase spectrum can be measured is limited to scales smaller than the average size of the dispersed phase. Moreover, the wavelet-based approach would also not generally work for compressible flows, because similar to a standard Fourier spectrum, the orthogonal wavelet spectrum does not satisfy the inviscid criterion. However, the spatial filtering approach we propose to compute spectra for two-phase flows in this work generally applies to incompressible/compressible two-phase flows and remedies the issue of the limit on the range of scales examined, as it seamlessly transitions from treating the dispersed phase as a union of disjoint regions to a full Eulerian field.
@@ 6.1 Computational setup
The final stationary single-phase HIT state at $t\approx 30\tau_{e}$ is used to initialize all the two-phase HIT simulations. This is done by inserting a sphere of another phase, $l=2$ (either a droplet or a bubble) in the center of the domain, as shown in Figure 3. To completely disregard the effect of the initialized velocity field inside the second phase (which could be unphysical), we run this setup using the various controlled forcings until the global surface area of the interface reaches a steady state, after which we collect statistics and make the findings reported below.
@@ 6.1.1 Initialization of the secondary phase: Effect of initial velocity field
Since the velocity field inside the second phase cannot be determined a priori, there have been different approaches to handling this velocity field inside the second phase in the past. Dodd & Ferrante (2016) zeroed out the velocity field when droplets of the second phase were introduced into a decaying isotropic turbulence, and Jain & Moin (2022) retained the same velocity as in the single-phase HIT when the slab and droplets of the second phase were introduced into an isotropic turbulence simulation.

### v2

@@ Abstract
keywords:
two-phase flows , turbulence , bubbles , droplets
journal: Chemical Engineering Journal
[…]
[label1]organization=Flow Physics and Computational Science Lab, George W. Woodruff School of Mechanical Engineering, Georgia Institute of Technology,state=GA, country=USA
[…]
[label2]organization=Center for Turbulence Research, Stanford University,state=CA, country=USA
@@ 1 Introduction
The deformation, fragmentation, and coalescence of bubbles and droplets in a turbulent environment is an important problem in various environmental and industrial settings (Liao and Lucas, 2009, 2010; Jain, 2017; Ni, 2024). In the atmosphere, the enhanced efficiency of water droplet collision–coalescence events due to air turbulence heavily influences cloud formation and warm rain initiation (Grabowski and Wang, 2013). At the ocean–atmosphere interface, underwater turbulence generated due to breaking waves breaks apart large entrained air pockets into smaller bubbles, facilitating mass transfer between the ocean and the atmosphere (Deike, 2022). Finally, turbulent emulsions, i.e., a homogenized mixture of two fluids of similar densities, are ubiquitously formed for use in pharmaceutical and industrial processing (Wang et al., 2007; Kilpatrick, 2012). As such, an understanding of how turbulence influences (and is influenced by) the dispersed phase, as parameters such as density and viscosity ratios, between the carrier and dispersed phases, as well as the volume fraction occupied by the dispersed phase are varied, is of paramount importance to understanding these ubiquitous flows.
@@
While these examples contain irreducible effects due to anisotropy, large-scale shear (Rosti et al., 2019), and buoyancy (Ravelet et al., 2011), the most straightforward setup to study the essential physics of deformation, fragmentation, and coalescence of bubbles and droplets in a turbulent environment is homogeneous isotropic turbulence (HIT). While simplistic, this setup allows for systematically studying the high-dimensional parameter space of density and viscosity ratios, volume fraction of dispersed phase, and Weber and Reynolds numbers. For each of these global parameters, the dependence of the phenomenology of the flow on the scales of turbulent fluctuations in the carrier phase and the scales of the dispersed phase can be further probed both above and below the Hinze scale (Hinze, 1955) (scale at which the inertial forces causing fragmentation due to turbulence are balanced by the restorative forces of surface tension). Incorporating the appropriate phenomenology of fragmentation and coalescence as a function of the global and scale-dependent parameters is essential to develop coarse-grained statistical models that can accurately capture the quantitative properties of the mixture. For example, while similar density mixtures lead to large-scale filamentary deformations requiring capillary-like breakup models (Eastwood et al., 2004; Kim and Moin, 2020), large density ratios such as in the case of air-water bubble break-up might be better modeled using a cascade-like model that may or may not account for the intermittent nature of turbulence (Chan et al., 2021a, b; Qi et al., 2022).
@@
Recent studies that utilize two-phase HIT as a testbed can be broken down into two categories: (i) The first category considers the detailed dynamics of individual bubbles and droplets, or emulsions, during the first few breakup events through simulations containing a single bubble/droplet, characterizing and quantifying things such as the generation of sub-Hinze scale bubbles due to rapid successive breakup events and capillary fragmentation (Riviere et al., 2021; Rivière et al., 2022), the effect of eddies far away from the two-phase interface on the deformation process (Vela-Martín and Avila, 2021), bubble deformation modes due to turbulent fluctuations (Perrard et al., 2021), the non-vanishing probability of breakup at sub-Hinze scales due to increasingly intermittent turbulent fluctuations (Vela-Martín and Avila, 2022), and the role of the viscosity ratio between the two phases on the rate of bubble breakup and the critical Weber number at which they break (Farsoiya et al., 2023). (ii) The second category considers characterizing the statistical properties of numerous bubbles and droplets present simultaneously in a turbulent environment. Within this category, Dodd and Ferrante (2016) studied the turbulent kinetic energy (TKE) budgets of droplets at various density ratios within decaying HIT. Crialesi-Esposito et al. (2022, 2023) studied the influence of viscosity ratios and volume fractions of the dispersed phase at unity density ratios in a forced HIT setting on the scale-by-scale energy budgets while linking the droplet-size distribution (DSD) to the local dissipation rate of turbulence, once again highlighting the importance of intermittency in the breakup dynamics. Finally, Begemann et al. (2022) and Krzeczek et al. (2023) utilize a linear forcing approach to study the segregation properties of emulsions, highlighting that even minor variations in the density ratio of the two phases from slightly below to slightly above unity shifts the DSD to have fewer super-Hinze bubbles and more sub-Hinze scale drops.
@@
Beyond this uncertainty of non-physical initial conditions, studies using decaying turbulence cannot separate the effect of the non-equilibrium change in the global properties of the surrounding carrier phase turbulence on the extracted statistics of the fragmentation processes for droplets/bubbles with timescales commensurate with some of the energetic scales of turbulence. In particular, for bubbles/droplets placed in the inertial range of turbulence, the recent evidence for the importance of a range of scales of smaller eddies on the break-up process could be drastically different if those small scales are decaying in time (Qi et al., 2022). Controlling this non-equilibrium effects would allow for more accurate measurements of the equilibrium statistics against which the non-equilibrium cases can be compared and interpreted as either slight deviations or completely different phenomena. As such, a statistically stationary state is sought instead. To achieve this state, forcing is necessary to sustain the continuously dissipating turbulence, ensuring that the turbulence properties do not change during the breakup processes. In the cases where forcing has been used, both simultaneous (both phases) and individual phase (carrier), linear, physical-space forcing (Lundgren, 2003; Rosales and Meneveau, 2005) have been utilized in various studies. However, individual phase (carrier) forcing is only used for the cases that study the break-up of single bubbles in HIT (Riviere et al., 2021; Rivière et al., 2022), and are hence more likely influenced by the choice of the initial condition as the simulations are only performed up to first few break-up events. In contrast, simultaneous mixture-phase (both phases) forcing has been applied only to density ratios representative of emulsions (Boukharfane et al., 2021; Begemann et al., 2022; Crialesi-Esposito et al., 2022; Krzeczek et al., 2023; Crialesi-Esposito et al., 2023; Boniou et al., 2025). As such, there remain open questions concerning the differing effects of simultaneous mixture-phase (both phases) forcing and individual phase (carrier) forcing, as well as how the aggregate properties of the two-phase HIT vary with density and viscosity ratios when forced.
@@ 2.1 Phase-field model
Note that a small number $\varepsilon$ is added to both the numerator and denominator to avoid division by 0, and $\varepsilon$ is chosen as $\varepsilon=10^{-100}$. The right-hand-side (RHS) term of Eq. (1) is the interface-regularization term, which is an artificial term that is added to maintain the constant thickness of the interface. The interface-regularization term contains a diffusion and a sharpening term, which balance each other to maintain a constant interface thickness on the order of $\epsilon$. The ACDI model is discretized using a second-order skew-symmetric split central scheme (Jain and Moin, 2022) that has superior numerical properties, such as improved robustness and lower aliasing error.
@@
This recently developed ACDI method has been shown to be suitable for complex turbulent flow simulations (Jain, 2022; Hwang and Jain, 2024) and a wide range of other multiphysics flow problems (Collis et al., 2022; Scapin et al., 2022; Brown et al., 2023; Jain, 2023, 2024). It is known to be more accurate than other existing phase-field models because it maintains a sharper interface—with only one-to-two grid points across the interface—while being non-dissipative, robust, less expensive, and conservative, without the need for any geometric treatment. Because of these advantages, the ACDI method is chosen as the interface-capturing method in this work.
@@ 2.3 Surface tension model
The surface tension forces, $f^{\sigma}_{i}=\sigma\kappa(\partial\phi/\partial x_{i})$ [last term in Eq. (3)], are modeled using a continuum-surface force (CSF) formulation (Brackbill et al., 1992), where $\sigma$ is the surface tension coefficient and $\kappa$ is the surface curvature, defined as
$\kappa=-(\partial n_{j}/\partial x_{j})=-\partial/\partial x_{j}\left[(\partial\psi/\partial x_{j})/\sqrt{(\partial\psi/\partial x_{i})(\partial\psi/\partial x_{i}})\right].$
The use of $\psi$, as opposed to $\phi$, to compute curvature in the surface tension force was shown in Jain (2022) to improve modeling of the surface tension force and reduce spurious currents significantly. This improved CSF model results in lower spurious currents compared to other existing surface tension models for diffuse-interface methods in the literature, including the free-energy based surface tension models (Jain, 2023). Hence, the improved CSF formulation is used to model surface tension forces in this work.
@@ 3.1 Mixture energy equations
where $k=\langle\rho u_{i}u_{i}/2\rangle$ is the mean mixture TKE, $\Psi=\langle u_{i}f_{i}^{\sigma}\rangle$ is the mean surface tension power, and $\epsilon=\langle\tau_{ij}{\partial u_{i}}/{\partial x_{j}}\rangle$ is the dissipation; for all these terms, the averaging operator is defined as $\langle\cdot\cdot\rangle=\left[\int_{\Omega}(\cdot\cdot)d\Omega\right]/\Omega$, where $\Omega$ is the domain volume. Similarly, we can write the mean surface energy equation (Dodd and Ferrante, 2016) as
@@ 3.2 Phase-averaged energy equations
where $k_{l}=\langle\phi_{l}\rho_{l}u_{i,l}u_{i,l}/2\rangle/\langle\phi_{l}\rangle$ is the phase-averaged TKE, $V_{l}=\langle\phi_{l}\left({\partial u_{i,l}\tau_{ij,l}}/{\partial x_{j}}\right)\rangle/\langle\phi_{l}\rangle$ is the phase-averaged viscous transport contribution, $P_{l}=\langle\phi_{l}\left({\partial u_{j,l}p_{l}}/{\partial x_{j}}\right)\rangle/\langle\phi_{l}\rangle$ is the phase-averaged pressure transport contribution, and $\epsilon_{l}=\langle\phi_{l}\tau_{ij}\left({\partial u_{i,l}}/{\partial x_{j}}\right)\rangle/\langle\phi_{l}\rangle$ is the dissipation of phase $l$. These equations are consistent with the diffuse-interface/phase-field formulation, where $\phi_{l}$ here represents the phase-field variable [see Dodd and Ferrante (2016) for a sharp-interface formulation of phase-averaged TKE equations].
@@ 5 Spatial filtering approach for two-phase spectra
To study the energy content at different scales of each of the dispersed and carrier phases in a two-phase flow, we compute the phase-dependent TKE spectra utilizing the spatial filtering approach of Sadek and Aluie (2018). Note that a standard Fourier spectrum is not a general enough approach for when there are variations in density and viscosity between the two phases. This is due to three reasons: First, the viscosity variations introduce discontinuities of velocity derivatives at interfaces, contaminating the spectrum. Second, using a Fourier spectrum, one cannot isolate the dispersed phase and carrier phase contributions. Third, and perhaps most importantly, energy is a cubic quantity in compressible flows and volumetrically filtered variable-density flows such as non-unity density ratio incompressible two-phase flows. However, energy must be defined quadratically using a Fourier spectrum; no longer satisfying the inviscid criterion (Aluie, 2013; Zhao and Aluie, 2018), i.e., the lack of influence of viscous effects at the large scales.
Using the wavelet method of Freund and Ferrante (2019), spectra have been computed before for the dispersed and carrier phases for some wavenumbers in an incompressible two-phase flow. However, due to the need for decomposing spatial regions into carrier, dispersed, and interaction phases, the range of scales over which a dispersed phase spectrum can be measured is limited to scales smaller than the average size of the dispersed phase. Moreover, the wavelet-based approach would also not generally work for compressible flows, because similar to a standard Fourier spectrum, the orthogonal wavelet spectrum does not satisfy the inviscid criterion. However, the spatial filtering approach we propose to compute spectra for two-phase flows in this work generally applies to incompressible/compressible two-phase flows and remedies the issue of the limit on the range of scales examined, as it seamlessly transitions from treating the dispersed phase as a union of disjoint regions to a full Eulerian field.
@@ 6.1 Computational setup
The final stationary single-phase HIT state at $t\approx 30\tau_{e}$ is used to initialize all the two-phase HIT simulations. This is done by inserting a sphere of another phase, $l=2$ (either a droplet or a bubble) in the center of the domain, as shown in Figure 3. To completely disregard the effect of the initialized velocity field inside the second phase (which could be unphysical), we run this setup using the various controlled forcings until the global surface area of the interface reaches a steady state, after which we collect statistics and make the findings reported below. The grid size of $N=256$ is chosen such that the global surface area is grid converged (Boukharfane et al., 2021; Jain and Elnahhas, 2022).
@@ 6.1.1 Initialization of the secondary phase: Effect of initial velocity field
Since the velocity field inside the second phase cannot be determined a priori, there have been different approaches to handling this velocity field inside the second phase in the past. Dodd and Ferrante (2016) zeroed out the velocity field when droplets of the second phase were introduced into a decaying isotropic turbulence, and Jain and Moin (2022) retained the same velocity as in the single-phase HIT when the slab and droplets of the second phase were introduced into an isotropic turbulence simulation.